# LocCa: Visual Pretraining with Location-aware Captioners

### v1

@@ Abstract
Image captioning has been shown as an effective pretraining method similar to contrastive pretraining. However, the incorporation of location-aware information into visual pretraining remains an area with limited research. In this paper, we propose a simple visual pretraining method with location-aware captioners (LocCa). LocCa uses a simple image captioner task interface, to teach a model to read out rich information, i.e. bounding box coordinates, and captions, conditioned on the image pixel input. Thanks to the multitask capabilities of an encoder-decoder architecture, we show that an image captioner can easily handle multiple tasks during pretraining. Our experiments demonstrate that LocCa outperforms standard captioners significantly on localization downstream tasks while maintaining comparable performance on holistic tasks.
Keywords:
Localization Image Captioning Vision Language Models
@@ 1 Introduction
In light of this, the modern contrastive pretraining methods [61, 31] extract learning signals from web-crawled image-text pairwise datasets [66, 12], circumventing the need for extensive manual annotations. The contrastively pretrained models demonstrate remarkable capability on zero-shot transfer tasks, especially on downstream applications that require fine-grained visual [90, 48] or textual understanding [98]. Besides, image captioning has been shown as an alternative visual pretraining task to learn capable vision encoders [72], where an encoder-decoder architecture is pretrained to generate text captions from image input. Some studies [93, 42] pioneered the cohesive pretraining of contrastive and generative methods. Typically, encoded image features are fed into two parallel branches: one employs a text encoder to produce sentence embeddings for contrastive learning, while the other utilizes a text decoder to generate image captions. Despite the effectiveness of these works, they typically focus on a holistic understanding of images, often overlooking the region-specific details of the visual content.
@@
To enhance the visual representation with location-aware context, [43, 100, 101] investigate the matching of image regions with corresponding text during pretraining. The central concept involves extracting Region of Interest (RoI) features from image embedding to facilitate contrastive learning with corresponding textual features. These approaches yield encouraging outcomes in location-sensitive tasks, such as object detection [23, 64, 45, 27] and referring expression [35, 95, 94, 60]. However, they necessitate complex model architectures (e.g., RPN [64] and FPN [45]) for RoI generation. Moreover, given the presence of multiple object candidates within a single image, region-wise matching becomes computationally demanding.
@@
Figure 1: Overview of LocCa. LocCa consists of a standard vision transformer and a transformer decoder. The vision transformer takes image pixel as input, produces visual tokens as cross attention input to the transformer decoder. The transformer decoder is trained to read out rich information from the visual tokens. We adopt the following three task for pretraining: (1) Captioning: generates a full caption for the given image; (2) Referring expression: generates a caption and the corresponding bounding box coordinates; (3) Grounded image captioning: generates a bounding box coordinates and the corresponding caption.
@@
Fortunately, the advancements in multitasking learning of decoders [3, 78, 50] have opened new pathways for integrating location information into visual representations. In this paper, we propose LocCa, a simple yet effective location-aware captioning pretraining method as shown in Figure 1, which uses an autoregressive decoder as an interface to handle additional location-aware pretraining tasks. Concretely, other than the image-to-text captioning task, LocCa also pretrains the model with two location-aware tasks: (i) referring expression from image and regional captions to predict bounding box coordinates, and (ii) grounded image captioning from image and bounding box coordinates to predict captions. Specifically, LocCa leverages a multi-task decoder [3] for pretraining, where the model outputs are conditioned on the task prefixes for each task. Thanks to the shared vision transformer for multiple tasks, the additional localization losses are relatively cheap, while the model inference speed is identical to the standard image caption pre-trained models.
@@
Our experimental results show that while maintaining the same level of capabilities on holistic tasks, the new LocCa performs significantly better on downstream tasks that require localization capabilities. We summarize our contributions as follows: (i) We propose a simple multi-task captioner that incorporates location information for visual pretraining; (ii) Without bells and whistles, LocCa achieves state-of-the-art results on localization tasks, while preserving the competitive performance on holistic tasks; and (iii) When integrated in vision-language models, i.e. PaLI-3 [11], the vision encoder outperforms strong SigLIP baselines [98].
@@ 2 Related Works
A natural alternative to contrastive pretraining is image captioning: Rather than matching image and text embeddings, one tries to predict n-grams in the caption or the full caption from an image embedding. Early works investigate this approach at small scale [33, 41, 17, 65]. Later works augment large-scale contrastive pretraining with a captioning loss [93, 42, 38], or scale-up captioning as a stand-alone task without investigating transfer to a broad set of vision tasks [29, 80]. [72] recently showed that image captioning alone leads to visual representations competitive with contrastive pretraining.
@@ 3 Location-aware Captioner
In this section, we introduce the location-aware image captioner LocCa for multitask pretraining. LocCa builds on an image captioner but provides a straightforward recipe for integrating location-aware information during model pretraining.
@@ 3.1 Pretraining tasks
The pretraining phase of LocCa draws inspiration from pioneering works that have successfully integrated a unified decoder for multitasking [78, 12, 50, 104, 85], utilizing a task-specific prefix for each distinct task. This methodology aligns with practices established by notable research, enhancing the model’s ability to handle multiple tasks concurrently.
For conventional image captioning, the process involves taking an image $x$ as input and generating a a sequence of text tokens $y=[y_{1},y_{2},\dots,y_{n}]$. In the LocCa framework, a task-specific prefix, labeled as “Cap:”, is added to the beginning of the caption sequence to designate the task at hand.
Moreover, LocCa integrates two additional location-aware tasks during its pretraining phase: referring expression (REF) [35, 95, 94, 60] and grounded captioning (GCAP) [32, 37, 81, 91, 20]. The former aims to pinpoint specific locations based on textual descriptions, while the latter focuses on creating captions that are directly tied to the precise location of objects within an image. These tasks are carefully chosen to ensure that LocCa achieves a fine-grained understanding of the intricate relationship between visual elements and textual semantics.
The foundation of LocCa’s pretraining is built upon dense regional annotations. Each image $x$ is associated with a comprehensive set of annotations $\{(b,c)\}$, where $b\in\mathcal{N}^{4}$ denotes the bounding box coordinates, and $c$ represents the corresponding textual descriptions or labels. For every bounding box identified, two distinct prompts are generated to cater to the aforementioned location-aware tasks: “Ref: $\{c\}$ : $\{b\}$” for referring expression and “GCap: $\{b\}$ : $\{c\}$” for grounded captioning, each prefixed with “Ref:” and “GCap:” respectively. These prompts are then tokenized to produce the sequence $y$ for each task, facilitating a targeted and effective pretraining regimen.
For each image, LocCa utilizes the identical visual feature extracted by the image encoder and executes three tasks using the same decoder in parallel. This pretraining scheme aims to make LocCa adept at linking fine-grained regional visual elements with appropriate textual descriptions, enhancing its capability to understand and generate related content effectively.
3.2 Model structure
LocCa utilizes a conventional encoder-decoder framework, where the encoder comprises a Vision Transformer [18] to transform the input image $x$ into a sequence of feature embeddings. The decoder, built on a Transformer architecture [73], processes these image features, employing cross-attention across each layer to facilitate the integration of visual and textual information effectively.
@@ 3.2 Model structure
3.3 Model training
The optimization of LocCa’s parameters $\theta$ is achieved through the maximization of the log-likelihood: $\sum_{i=1}^{|y|}\log P_{\theta}(y_{i}|y_{<i},x)$. It is important to emphasize that, in the learning process for the tasks of referring expression and grounded captioning, LocCa is structured to predict captions and bounding boxes sequentially, contrasting with traditional approaches that might predict a caption based on a given bounding box or vice versa.
This is achieved by applying losses to the entire prompt excluding the task prefix. Taking the referring expression task as an example, the overall loss is computed for both textual predictions $c$ and box coordinates $b$. The loss on $c$ guides the model to identify a foreground region and generate its caption, while the loss on $b$ aims at refining the model’s ability to accurately regress the box location relative to the caption. This dual-faceted loss structure ensures that the generated caption and the corresponding bounding box are contextually and spatially coherent, enhancing the model’s understanding of the visual and textual elements in unison.
Moreover, this design enhances inference flexibility of LocCa, allowing for varied input configurations. Users can input a single task prefix (e.g., “Ref:”) to prompt the model to identify and describe an area of interest along with its location. Alternatively, by inputting both the task and a conditional input (e.g., “Ref: a black and white cat :”), LocCa can be directed to focus solely on predicting the location. This flexibility allows for customized responses to various inquiries, highlighting the model’s adaptability to meet specific user needs.
@@ 4.1 Experimental setup
We use a subset of the WebLI dataset [12] corresponding to English websites and apply text-based filtering [31] to obtain 1B image/alt-text pairs. It is worth noting that WebLI data was de-duplicated w.r.t. all the images in the evaluation data sets used in this paper. To obtain fine-grained object locations, a publicly available OWL-ViT CLIP L/14 model [53] is applied to generate detection pseudo annotations. Specifically, two groups of box categories are generated: the n-gram texts from alt-text and the object categories as used by PaLI [12], more details can be found in [52]. In LocCa pretraining, we first filter the candidate bounding boxes according to their OWL-ViT confidence score with a threshold of 0.3, and then randomly sample one box-caption or box-category pair for referring expression and grounded captioning separately.
@@
Model details
LocCa adopts an encoder-decoder structure where, unless otherwise noted, the encoder defaults to a standard ViT-L/14 design with 24 transformer blocks handling input image patches of size 14. The decoder, following [72], is a Transformer-L model consisting of 12 transformer decoder blocks. In total, the model comprises approximately 0.6B parameters. Two more LocCa setups with ViT-B/16 and ViT-G/14 (see [97] for model specifications) are adopted for ablation tests and scaling experiments respectively.
@@
LocCa is pretrained for about 9 billion image/alt-text seen examples, which corresponds to about 9 epochs on our tailored subset of WebLI. For the optimizer, we employ the Scaling-ViT AdaFactor variant [97], combined with a cosine schedule that includes 10,000 warmup steps. The batch size is set at 8,192, while the learning rate and decay factor are adjusted to $10^{-3}$ and $10^{-4}$, respectively. During this process, images are uniformly resized to a resolution of 224 x 224 pixels. Alt-texts are tokenized into a vocabulary consisting of 32,000 tokens using a sentence piece model trained on the English segment of C4 [62], with a cap on the sequence length at 64 tokens. We represent bounding box coordinates using up to 500 integral numbers, which are then directly converted into strings for straightforward representation of coordinate tokens. For parallel prediction in the vanilla image captioning task, the fraction of examples is set to 50% by default.
Selection of downstream tasks
Our goal with LocCa is to preserve or even improve the capability on various image-level understanding tasks, and get a higher performance on fine-grained location-aware tasks. To this end, we assess the capabilities of LocCa in holistic and location-aware image understanding tasks across a range of downstream tasks. In the realm of holistic image understanding, we focus on the same LiT-Decoder tasks [3] in CapPa [72] including image classification (CLS)[16, 7, 83, 15, 58], image captioning (CAP) [13, 92], optical character recognition (OCR-VQA) [54], visual question answering (VQA) [24], and graph question answering (GQA) [30]. For evaluating location-aware image understanding, we choose two widely recognized tasks like referring expression comprehension (REC) [35], referring expression segmentation (RES) [28] and object detection (OD) [64]. Additionally, paralleling the approaches of PaLI [12, 11], we integrate LocCa to a pretrained large language model to assess performance on a variety of vision-language tasks, including image captioning and visual question answering.
Notably, despite LocCa not being exposed to any video content during pretraining, we adapt it to various video-related tasks, such as video captioning, QA, and classification. This adaptation aims to assess its generalization capabilities to new modalities, demonstrating its versatility compared to other image-text pretraining approaches.
Transfer to downstream tasks
@@ 4.2 Quantitative results
We conduct extensive experiments to validate the advancement of LocCa. The integration of location-aware cues enables LocCa to maintain its performance on holistic image understanding tasks while achieving substantially improved outcomes on location-aware tasks. Further enhanced by an advanced pretrained large language model, LocCa exhibits exceptional performance across a range of vision-language tasks, substantially surpassing baseline models.
@@ 4.2.1 Referring Expression Comprehension
In this section, we present results on the referring expression comprehension benchmarks, including RefCOCO, RefCOCO+ and RefCOCOg [95]. As shown in Table 1, LocCa establishes a new state-of-the-art across these benchmarks thanks to the location-aware pretraining. This advancement is particularly noteworthy considering the inherent limitations of previous methods. For example, UNINEXT[88], which adopts a Deformable DETR architecture[103] tailored for detection-based tasks, and OFA, which requires a pretrained BART[40] for initialization, both achieve respectable results but are constrained by their specialized setups. In contrast, LocCa employs a standard encoder-decoder architecture with auto-regressive pretraining from scratch, significantly outperforming these methods across all benchmarks without the need for complex task-specific adaptations.
Achieving good performance on RefCOCO usually requires pretraining at high image resolutions. For instance, OFAL is pretrained at 4802px, while UNI-NEXT undergoes multi-scale pre-training. We opt for a standard 2242px pretrain resolution for simplicity. We transfer the 2242px pretrain model to RefCOCO by fine-tuning with 6402px resolution, using learning rate $10^{-4}$ and no weight decay. We report the standard metric Acc@0.5 on the validation and test sets.
Notably, we train on the combined training sets of RefCOCO, RefCOCO+ and RefCOCOg but removing all validation and test images from this combination. The splits of these three datasets are largely overlapping, meaning that methods trained on the combined training sets without de-duplication, a recently common phenomenon, have trained on about half of the test images, see Appendix 0.A.2 for details. We provide the list of removed image IDs in the Appendix 0.A.2. Furthermore, some models such as UNITER [14], VILLA [22], and MDETR [34] use COCO pre-trained detection components which have seen test images from the three RefCOCO versions. We have removed all training, validation, and test images from the COCO dataset from our pre-training data, as well as near-duplicates thereof. Hence, we group models reported in Table 1 into three distinct categories. We transfer LocCa on both the full and the “clean” combined training set and provide both results as a point of comparison. We hope that, going forward, the community evaluates models on RefCOCO in this clean setup.
Table 2: Comparison with baselines on RefCOCOs. A randomly initialized decoder is trained for REC and RES, keeping the image representation frozen. We report Acc@0.5 for REC and mIoU for RES. Here CLIP uses model checkpoints released by [61]; all other baselines are trained on the same data as LocCa.

### v2

@@ Abstract
Image captioning was recently found to be an effective pretraining method similar to contrastive pretraining. This opens up the largely-unexplored potential of using natural language as a flexible and powerful interface for handling diverse pretraining tasks. In this paper, we demonstrate this with a novel visual pretraining paradigm, LocCa, that incorporates location-aware tasks into captioners to teach models to extract rich information from images. Specifically, LocCa employs two tasks, bounding box prediction and location-dependent captioning, conditioned on the image pixel input. Thanks to the multitask capabilities of an encoder-decoder architecture, we show that an image captioner can effortlessly handle multiple tasks during pretraining. LocCa significantly outperforms standard captioners on downstream localization tasks, achieving state-of-the-art results on RefCOCO/+/g, while maintaining comparable performance on holistic tasks. Our work paves the way for further exploration of natural language interfaces in visual pretraining.
@@ 1 Introduction
In light of this, the modern contrastive pretraining methods [7, 8] extract learning signals from web-crawled image-text pairwise datasets [9, 10], circumventing the need for extensive manual annotations. The contrastively pretrained models demonstrate remarkable capability on zero-shot transfer tasks, especially on downstream applications that require fine-grained visual [11, 12] or textual understanding [13]. More recently, image captioning has been shown as an alternative visual pretraining task to learn capable vision encoders [14], where an encoder-decoder architecture is pretrained to generate text captions from the image input. Some studies, such as [15, 16], pioneered the joint pretraining of contrastive and generative methods. Typically, encoded image features are fed into two parallel branches: one employs a text encoder to produce sentence embeddings for contrastive learning, while the other utilizes a text decoder to generate image captions. Despite the effectiveness of these works, they typically focus on a holistic understanding of images, often overlooking the region-specific details of the visual content.
@@
The recent success of image captioning [14] and the advancements in multitasking learning of decoders [17, 18, 19] opens up the largely-unexplored potential of using natural language as a flexible and powerful interface for handling diverse tasks. We demonstrate this with a novel visual pretraining paradigm, LocCa, that enhances the visual representation with location-aware context. Works including [20, 21, 22] investigate the matching of image regions with corresponding text during pretraining. The central concept involves extracting Region of Interest (RoI) features from image embedding to facilitate contrastive learning with corresponding textual features. These approaches yield encouraging outcomes in location-sensitive tasks, such as object detection [23, 24, 25, 26] and referring expression [27, 28, 29, 30]. However, they require complex model architectures (e.g. RPN [24] and FPN [25]) for RoI generation. Also, given the presence of multiple object candidates within an image, region-wise matching becomes computationally demanding.
@@
Figure 1: Overview of LocCa. LocCa consists of a standard vision transformer and a transformer decoder. The vision transformer takes image pixel as input, produces visual tokens as cross attention input to the transformer decoder. The transformer decoder is trained to read out rich information from the visual tokens. We adopt the following three task for pretraining: Cap, AREF and GCAP.
@@
By contrast, LocCa is a simple yet effective location-aware captioning pretraining method as shown in Figure 1, which uses an autoregressive decoder as an interface to handle additional location-aware pretraining tasks. Concretely, other than the image-to-text captioning task, LocCa also pretrains the model with two location-aware tasks: (i) automatic referring expressions, which amounts to predict bounding box coordinates from automatically generated captions for specific image regions, and (ii) grounded captioning to jointly predict box coordinates and captions from the image. Specifically, LocCa leverages a multi-task decoder [17] for pretraining, where the model outputs are conditioned on the task prefixes for each task. Thanks to the shared vision transformer for multiple tasks, the additional localization losses are relatively cheap to compute, while the model inference speed is identical to the standard image caption pretrained models.
@@
Our experimental results show that the LocCa performs significantly better on downstream tasks that require localization capabilities, while maintaining the same level of capabilities on holistic tasks. We summarize our contributions as follows: (i) For the first time we explore location-aware tasks as proxies for generative visual pretraining (as opposed to transfer/instruction tuning in prior works), enabling flexibly customized inference (detailed in Sec.3.3); (ii) Without bells and whistles, LocCa achieves state-of-the-art results on localization tasks, while preserving the competitive performance on holistic tasks; and (iii) When integrated in vision-language models, i.e. PaLI-3 [31], the vision encoder outperforms strong SigLIP baselines [13].
@@ 2 Related Works
A natural alternative to contrastive pretraining is image captioning: Rather than matching image and text embeddings, one tries to predict captions from an image embedding. Early works investigate this approach at small scale [38, 39, 40, 41]. Later works augment large-scale contrastive pretraining with a captioning loss [15, 16, 42], or scale-up captioning as a stand-alone task without investigating transfer to a broad set of vision tasks [43, 44]. [14] recently showed that image captioning alone leads to visual representations competitive with contrastive pretraining.
@@ 3 Location-aware Captioner
In this section, we introduce the location-aware image captioner LocCa for multitask pretraining. LocCa builds on an image captioner but provides a recipe for integrating location-aware information during model pretraining.
@@ 3.1 Pretraining tasks
The pretraining phase of LocCa draws inspiration from pioneering works that have successfully integrated a unified decoder for multitasking based on pretrained models [18, 10, 19, 46, 52], utilizing a task-specific prefix for each distinct task. This enhances the model’s ability to handle multiple tasks concurrently.
For conventional image captioning, the process involves taking an image $x$ as input and generating a sequence of text tokens $y=[y_{1},y_{2},\dots,y_{n}]$. In the LocCa framework, a task-specific prefix, labeled as “Cap:”, is added to the beginning of the caption sequence to designate the task at hand. Moreover, LocCa integrates two additional location-aware tasks during its pretraining phase: automatic referring expression (AREF) and grounded captioning (GCAP). These tasks are inspired by referring expression comprehension [27, 28, 29, 30] and dense captioning [56, 57, 58, 59, 60] respectively. The key difference is that LocCa predicts both regional captions and box coordinates sequentially with task prefixes, instead of relying on either caption or box conditional inputs (see Fig. 1).
The foundation of LocCa’s pretraining is built upon dense, automatically generated region annotations. Each image $x$ is associated with a comprehensive set of annotations $\{(b,c)\}$, where $b\in\mathcal{N}^{4}$ denotes the bounding box coordinates, and $c$ represents the corresponding textual descriptions or labels. For every bounding box, two distinct prompts are generated to cater to the aforementioned location-aware tasks: “ARef: $\{c\}$ : $\{b\}$” for automatic referring expression and “GCap: $\{b\}$ : $\{c\}$” for grounded captioning, each prefixed with “ARef:” and “GCap:” respectively. These prompts are then tokenized to produce the sequence $y$ for each task, facilitating pretraining with a text interface.
For each image, LocCa utilizes the same visual features extracted by the image encoder and performs three tasks using the same decoder in parallel. This pretraining scheme aims to make LocCa adept at linking fine-grained regional visual elements with appropriate textual descriptions.
3.2 Model details
Architecture
LocCa utilizes a conventional encoder-decoder framework, where the encoder comprises a Vision Transformer [2] to transform the input image $x$ into a sequence of feature embeddings. The decoder, built on a Transformer architecture [61], processes these image features, employing cross-attention across each layer to integrate visual and textual information effectively.
Autogressive decoding
@@ 3.2 Model details
Objective
The optimization of LocCa’s parameters $\theta$ is achieved through the maximization of the log-likelihood: $\sum_{i=1}^{|y|}\log P_{\theta}(y_{i}|y_{<i},x)$. It is important to emphasize that, in the learning process for the location-aware tasks, LocCa is structured to predict captions and bounding boxes sequentially, contrasting with traditional approaches that might predict a caption based on a given bounding box or vice versa [18, 51]. This is achieved by applying losses to the entire prompt excluding the task prefix. Taking the AREF task as an example, the overall loss is computed for both textual predictions $c$ and box coordinates $b$. The loss on $c$ guides the model to identify a foreground region and generate its caption, while the loss on $b$ aims at refining the model’s ability to accurately regress the box location relative to the caption.
3.3 Discussion
To the best of our knowledge, LocCa is the first end-to-end method that incorporates the location-aware tasks (i.e., AREF and GCAP) into generative VLM pretraining. Our key novelty lies in the formulation of the location-aware proxy tasks which allows for scalable pretraining and enhances the visual localization capabilities. Compared to [22, 18, 51, 52] that require either a pretrained visual encoder or language decoder, LocCa does not need any pretrained model for initialization. Compared to [22, 62] that employ complex architectures with multiple losses, LocCa adopts a simple encoder-decoder architecture with a single generative loss. Compared to [18, 19] that are pretrained on a large number of tasks, LocCa introduces the novel use of simple AREF and GCAP proxy tasks for visual pretraining.
The dual-faceted loss structure of AREF and GCAP achieves comparable results to directly adopting the traditional referring expression and dense captioning tasks. However, it enhances inference flexibility of LocCa, allowing for varied input configurations. For instance, users can input a single task prefix (e.g., “ARef:”) to prompt the model to identify and describe an area of interest along with its location. Alternatively, by inputting both the task and a conditional input (e.g., “ARef: a black and white cat :”), LocCa can be directed to focus solely on predicting the location. This flexibility allows for customized responses to various inquiries, highlighting the model’s adaptability to meet specific user needs.
@@ 4.1 Experimental setup
We use a subset of the WebLI dataset [10] corresponding to English websites and apply text-based filtering [8] to obtain 1B image/alt-text pairs. The WebLI data was de-duplicated w.r.t. all the images in the evaluation data sets used in this paper. To obtain fine-grained object locations, a publicly available OWL-ViT CLIP L/14 model [63] is applied to generate detection pseudo annotations. Specifically, two groups of box categories are generated: the n-gram texts from alt-text and the object categories as used by PaLI [10], more details can be found in [64]. In LocCa pretraining, we first filter the candidate bounding boxes according to their OWL-ViT confidence score with a threshold of 0.3, and then randomly sample one box-caption or box-category pair for referring expression and grounded captioning separately.
@@
Implementation details
LocCa adopts an encoder-decoder structure where, unless otherwise noted, the encoder defaults to a standard ViT-L/14 design with 24 transformer blocks handling input image patches of size 14. The decoder, following [14], is a Transformer-L model consisting of 12 transformer decoder blocks. In total, the model comprises approximately 600M parameters. Two more LocCa setups with ViT-B/16 and ViT-G/14 (see [4] for model specifications) are adopted for ablation tests and scaling experiments respectively.
@@
LocCa is pretrained for about 9 billion image/alt-text seen examples, which corresponds to about 9 epochs on our tailored subset of WebLI. For the optimizer, we employ the Scaling-ViT AdaFactor variant [4], combined with a cosine schedule that includes 10,000 warmup steps. The batch size is set at 8,192, while the learning rate and decay factor are adjusted to $10^{-3}$ and $10^{-4}$, respectively. During this process, images are uniformly resized to a resolution of 224 x 224 pixels. Alt-texts are tokenized into a vocabulary consisting of 32,000 tokens using a sentence piece model trained on the English segment of C4 [65], with a cap on the sequence length at 64 tokens. We represent bounding box coordinates using up to 500 integral numbers, which are then directly converted into strings for straightforward representation of coordinate tokens. For parallel prediction in the vanilla image captioning task, the fraction of examples is set to 50% by default. The pretraining of LocCaL takes 153 hours using 256 TPUv3 chips.
Downstream tasks
Our goal with LocCa is to preserve or even improve the capability on various image-level understanding tasks, and get a higher performance on fine-grained location-aware tasks. To this end, we assess the capabilities of LocCa in holistic and location-aware image understanding tasks across a range of downstream tasks. In the realm of holistic image understanding, we focus on the same LiT-Decoder tasks [17] in CapPa [14] including image classification (CLS)[5, 66, 67, 68, 69], image captioning (CAP) [70, 71], optical character recognition (OCR-VQA) [72], visual question answering (VQA) [73], and graph question answering (GQA) [74]. For evaluating location-aware image understanding, we choose two widely recognized tasks like referring expression comprehension (REC) [27], referring expression segmentation (RES) [75] and object detection (OD) [24]. Additionally, paralleling the approaches of PaLI [10, 31], we integrate LocCa to a pretrained large language model to assess performance on a variety of vision-language tasks, including image captioning and visual question answering. Notably, despite LocCa not being exposed to any video content during pretraining, we adapt it to various video-related tasks, such as video captioning, QA, and classification. This adaptation aims to assess its generalization capabilities to new modalities, demonstrating its versatility compared to other image-text pretraining approaches. We adopt different strategies for transferring LocCa to downstream tasks, as summarized in Appendix A.1.
@@ 4.2 Quantitative results
We conduct extensive experiments to evaluate LocCa. The integration of location-aware cues enables LocCa to maintain its performance on holistic image understanding tasks while achieving substantially improved outcomes on location-aware tasks. Further enhanced by an advanced pretrained large language model, LocCa exhibits exceptional performance across a range of vision-language tasks, substantially surpassing baseline models.
@@ 4.2.1 Referring Expression Comprehension
In this section, we present results on the referring expression comprehension benchmarks, including RefCOCO, RefCOCO+ and RefCOCOg [28]. As shown in Table 1, LocCa establishes a new state-of-the-art across these benchmarks. This advancement is particularly noteworthy considering the inherent limitations of previous methods. For example, UNINEXT [77], which adopts a Deformable DETR architecture [84] tailored for detection-based tasks, and OFA, which requires a pretrained BART [85] for initialization, both achieve good results but are constrained by their specialized setups. Besides, some location-aware VLMs, such as Shikra [78] and Ferret [79], require an LLM for knowledge-based reasoning, which increases inference costs. In contrast, LocCa employs a standard encoder-decoder architecture with auto-regressive pretraining from scratch, significantly outperforming these methods across all benchmarks without the need for complex task-specific adaptations.
Achieving good performance on RefCOCO usually requires pretraining at high image resolutions. For instance, OFAL is pretrained at 4802px, while UNI-NEXT undergoes multi-scale pre-training. We opt for a standard 2242px pretrain resolution for simplicity. We transfer the 2242px pretrained model to RefCOCO by fine-tuning with 6402px resolution, using learning rate $10^{-4}$ and no weight decay. We report the standard metric Acc@0.5 on the validation and test sets.
Notably, we train on the combined training sets of RefCOCO, RefCOCO+ and RefCOCOg but removing all validation and test images from this combination. The splits of these three datasets are largely overlapping, meaning that methods trained on the combined training sets without de-duplication, a recently common phenomenon, have trained on about half of the test images, see Appendix A.3 for details. We provide the list of removed image IDs in the Appendix A.3. Furthermore, some models such as UNITER [80], VILLA [81], and MDETR [82] use COCO pre-trained detection components which have seen test images from the three RefCOCO versions. We have removed all training, validation, and test images from the COCO dataset from our pre-training data, as well as near-duplicates thereof. Hence, we group models reported in Table 1 into three distinct categories. We transfer LocCa on both the full and the “clean” combined training set and provide both results. We hope that, going forward, the community evaluates models on RefCOCO in this clean setup.
Table 2: Comparison with baselines on RefCOCOs. A randomly initialized decoder is trained for REC and RES, with frozen image encoders. We report Acc@0.5 for REC and mIoU for RES. Here CLIP uses model checkpoints released by [7]; all other baselines use the same data as LocCa.